The Company is in the process of adjusting and correcting its books at the end of...

  1. The Company is in the process of adjusting and correcting its books at the end of 2017. In reviewing its records, the following information is compiled.
  1. At December 31, 2017, the company decided to change the depreciation method on its office equipment from double-declining-balance to straight-line. The equipment had an original cost of $200,000 when purchased on January 1, 2015. It has a 10-year useful life and no salvage value. Depreciation expense recorded prior to 2017 under the double-declining-balance method was $72,000.   They have already recorded 2017 depreciation expense of $25,600 using the double-declining-balance method. Prepare the necessary journal entry at December 31, 2017 before they close their books.

Debit

Credit

  1. Insurance for a 12-month period purchased on November 1 of this year was charged to insurance expense in the amount of $3,300 because “the amount of the check is about the same every year.”   Prepare the necessary journal entry at December 31, 2017 before they close their books.

Debit

Credit

  1. Reported sales revenue for the year is $1,908,000. This includes all sales taxes collected for the year. The sales tax rate is 6%. Because the sales tax is forwarded to the state’s Department of Revenue, the Sales Tax Expense account is debited. The bookkeeper thought that “the sales tax is a selling expense.” At the end of the current year, the balance in the Sales Tax Expense account is $103,400.   Prepare the necessary journal entry at December 31, 2017 before they close their books.

Debit

Credit
Answer #1

(a. )

Change in method of depreciation is a change in accounting estimate.So change can be done prospectively .To change the method we need to know the Carrying amount of the asset on the date of change in method

In this case the Company had purchased asset in Jan 2015 with estimated cost of $ 200,000 with a life of 10 years with no salvage value. Depreciation till 2017 (for 2 years) is $ 72000 under double declining method.

Change happened in 2017 after 2 years

Carrying amount on the date of change is $200000-$72000=$128000

Remaining useful life after depreciating for 2 years under diminishing balance method is 8 years

So depreciation henceforth every year under straight line method for remaining years is $ 128000/8=$16000

In this case $25,600 has been already charged using double declining balance forthe year 2017 which is wrong as they have already adopted the change in the beginning of 2017

So difference in both the method =$ 25600-$16000=$9600 has to be rectified by increasing the value of asset for reducing the effect of depreciation over charged

So on Dec 31st 2017 Rectifying entry will be

Office Equipment $ 9600(Debit)

Depreciation expense $ 9600(Credit)

(Depreciation over charged has been reduced by crediting depreciation and the difference is added back to the Value of asset)

(b.)

Insurance for a 12 Month Period paid on Nov 1 is $ 3300. It includes Prepaid Insurance ..That amount which is paid in Nov that pertains to next year (Jan-Oct)is treated as prepaid and the amount paid for 2 months (NOv and Dec) is considered as Insurance Expenses for the Year

$3300*2/12=550 is Insurance Expense for the year(2 months-NOV &Dec)

$3300*10/12=2750 is Prepaid Insurance (paid in advance for next year which is considered as asset) so to reduce the effect and transferring the excess insurance charged to prepaid,

Rectifying Entry will be

Prepaid Insurance $ 2750(Debit)

Insurance Expense $ 2750(Credi)

(Insurance expenses is split between years and correspondingly reduced to record the difference as prepaid Insurance)

(c)Sales revenue that Included Sales tax of 6% can be Split into $ 108000 of sales tax and $ 1800000 as sales revenue

$ 1908000*6/106=$108000 and

$ 1908000*100/106=$ 1800000

In this case there are 2 errors

1.Sales Tax is reported Low so the difference of $ 108000 and $ 103400 is $4600 which needs to be shown as Sales Tax liability and paid to State's Department of Revenue

2.Sales Tax is not a Selling Expense.It is a Liability'(Error in Classification)

Actual entry Should have been (assuming Sales is realised in cash or if not it can be debited as Accounts Receivable)

Correct Entry:

Cash $ 1908000(Debit)

To Sales $ 1800000(Credit,being Revenue)

To Sales Tax Payable $ 108000(Credit,being Liability owed)

Entry that was passed

Cash $1908000

To Sales $ 1908000

and eventually the sales tax of $ 103400 was charged as sales tax expense and its payment entry is

Sales Tax Expense 103400(Debit)

Cash 103400(Credit)

So there are 2 errors one in classifying and other in understatement of Sales Tax amount for $ 4600

in this case there was a shortage of $ 4600 ($108000-$103400)in case of Sales tax so assuming it is included in sales ledger Account the rectifying entry will be

Sales $ 4600

Sales tax Payable $ 4600

Sales Tax is not a Selling expense it is a liability. So whole of sales tax expense already included in sales should be debited to report actual sales

Sales $103400 (Debit) because it was Previously credited)

Sales Tax Expense $ 103400

Alternatively combined entry can be

Sales $108000

To Sales Tax Payable $ 4600

To Sales Tax Expense $ 103400

So now by debiting sales A/c (4600+103400)=$108000 the sales is reduced to the extent of taxes which will now have a balance of $ 1800000 and Sales Tax expense will show nil balance and Difference of $ 4600 in taxes paid will be shown as Sales Tax Payable